OpenStudy (anonymous):

Which function has a domain of all real numbers? A. Y=tanx B. Y=cosx C. Y=1/sinx D. Y= secx

OpenStudy (sweetburger):

is there a difference between B. and C. that I am not seeing?

OpenStudy (anonymous):

I meant Y=1/sinx

OpenStudy (sweetburger):

ok I am pretty sure the domain of y=cosx continues into infinity in both directions+,- only the range is restricted from 1 to -1

OpenStudy (sweetburger):

y= tanx has aymptotes. I think y = secx has asymptotes and I also think that y= 1/sinx has asymptotes as well but Im a bit lazy to look them up.

OpenStudy (anonymous):

Think of it this way:\[A. y=\tan x = \frac{ \sin x }{ \cos x }\]\[B. y=\cos x\]\[C. y=\frac{ 1 }{ \sin x }\]\[D. y=\sec x = \frac{ 1 }{ \cos x }\]In order for the domain to be the set of all real numbers, there must not exist a real number that would make any denominator equal to zero, as that would lead to an undefined value. Does that help?

OpenStudy (anonymous):

So it would be B, y=cosx?

OpenStudy (anonymous):

Very good.